Verify a – (– b) = a + b for the following values of a and b.
a = 21, b = 18

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

From the question,
a = 21 and b = 18
To verify a–(- b) = a + b
Let us take Left Hand Side (LHS) = a–(- b)
= 21–(- 18)
= 21 + 18
= 39
Now, Right Hand Side (RHS) = a + b
= 21 + 18
= 39
By comparing LHS and RHS
LHS = RHS
39 = 39
Hence, the value of a and b is verified.

Thanks